# End-to-End Visual Autonomous Parking via Control-Aided Attention

### v1

@@ Abstract
Precise parking requires an end-to-end system where perception adaptively provides policy-relevant details—especially in critical areas where fine control decisions are essential. End-to-end learning offers a unified framework by directly mapping sensor inputs to control actions, but existing approaches lack effective synergy between perception and control. We find that transformer-based self-attention, when used alone, tends to produce unstable and temporally inconsistent spatial attention, which undermines the reliability of downstream policy decisions over time. Instead, we propose CAA-Policy, an end-to-end imitation learning system that allowes control signol to guide the learning of visual attention via a novel Control-Aided Attention (CAA) mechanism. For the first time, we train such an attention module in a self-supervised manner, using backpropagated gradients from the control outputs instead of from the training loss. This strategy encourages the attention to focus on visual features that induce high variance in action outputs, rather than merely minimizing the training loss—a shift we demonstrate leads to a more robust and generalizable policy. To further enhance stability, CAA-Policy integrates short-horizon waypoint prediction as an auxiliary task, and introduces a separately trained motion prediction module to robustly track the target spot over time. Extensive experiments in the CARLA simulator show that CAA-Policy consistently surpasses both the end-to-end learning baseline and the modular BEV segmentation + hybrid A* pipeline, achieving superior accuracy, robustness, and interpretability. Code is released at https://github.com/Joechencc/CAAPolicy.
@@ I Introduction
Autonomous driving has traditionally been approached with modular pipelines that decompose the task into perception, prediction, and planning/control [1, 2, 3]. While such designs provide interpretability and modularity [4, 5, 6], they are prone to error accumulation across stages [7], redundant computation [6, 8], and reduced robustness under perturbations [9],
Fig. 1LABEL:fig:teasing(a). End-to-end (E2E) learning offers a unified alternative by mapping sensory inputs directly to control outputs through a single neural network [10, 11, 12], Fig. 1LABEL:fig:teasing(b).
Existing end-to-end (E2E) autonomous parking methods [13] typically predict control commands directly from multi-view images. While these approaches unify perception and control, their synergy remains weak: perception modules are trained solely through gradients backpropagated from task losses, and thus often fail to focus on control-relevant cues such as parking slot boundaries, lane markings, and obstacle gaps. This limitation is particularly pronounced in attention mechanisms. As shown in Fig. 2(a–b), a baseline without control guidance scatters attention across irrelevant regions and exhibits temporal inconsistency. Moreover, even when incorporating a automatic learned attention module such as CBAM[14], the model still fails to consistently emphasize near-target areas. Without stronger coupling between perception and control, attention tends to drift toward spurious features, resulting in misaligned perception and unreliable policy outputs.
To address these limitations, we propose CAA-Policy, an end-to-end autonomous driving framework that tightly couples perception and control (Fig. 1LABEL:fig:teasing(c)). At its core, it introduces a self-supervised Control-Aided Attention (CAA) mechanism that learns attention from control-output gradients rather than the loss gradients, enabling the attention module consistently focuses on control-sensitive regions near the target slot (Fig. 2(c)). Driving goals are further encoded as structured, interpretable tokens for goal-aware control, and short-horizon waypoint prediction is added as an auxiliary task to improve temporal consistency. Together, these components yield more accurate and robust autonomous parking.
Figure 2: Attention maps for target tracking at two consecutive frames ($t$ and t+1t{+}1): (a) without CAA, attention scatters over irrelevant areas; (b) with modified CBAM [14], an automatically learned attention module from task supervision (no auxiliary loss); (c) with CAA, attention focuses on the target and its vicinity; (d) ground-truth segmentation (yellow: target, blue: obstacles).
@@
  • CAA-Policy, an end-to-end autonomous parking framework that tightly couples perception and control.
  • A novel Control-Aided Attention (CAA) mechanism that learn the attention model from control gradients wrt. perception features, making it consistantly focus on control-sensitive areas.
  • A learnable motion prediction module that tracks the vehicle’s current state using multiple historical motion frames for reliable current-step prediction.
  • A short-horizon waypoint prediction auxiliary task to enhance trajectory consistency and planning stability.
  • Extensive experiments demonstrate that CAA-Policy  significantly outperforms both modular hybrid A* and previous end-to-end baseline in trajectory accuracy, parking success rate, and interpretability in CARLA.
@@ II-A Autonomous parking system
Prior work on autonomous parking ranges from traditional modular pipelines to modern End-to-End (E2E) methods. Early systems [15, 16, 17] typically combined occupancy mapping, hand-crafted planners (e.g., hybrid A* [18], RRT [19]), and rule-based controllers, offering interpretability but limited adaptability. Recent advances largely adopt either modular learning approaches or E2E learning paradigms [20, 13, 21].
Modular approaches follow the classic decomposition of perception, planning, and control, with each stage designed and tuned separately. Modular pipelines with hybrid A* planners [22, 23] face fundamental efficiency limitations that directly impact accuracy under practical trade-offs. Owing to the breadth-first-search nature of hybrid A*, planning becomes increasingly costly when the target is far, as the frontier space grows with each step. To mitigate the high computational burden, systems often reduce the frequency of re-planning with new perception inputs. However, this leads to blind execution based on outdated information, where occlusions near the target are ignored, frequently resulting in failure. This shortcoming has motivated the subsequent preference for learning-based policies and end-to-end approaches.
End-to-end approaches provide a unified alternative to modular pipelines by directly mapping sensory inputs to control outputs, making them particularly suitable for structured tasks like autonomous parking. Within this paradigm, one line of work formulates parking as a trajectory prediction task, where CNN–LSTM or Transformer architectures are supervised with trajectory waypoints rather than direct control commands [24, 25, 21]. Another line of work adopts imitation learning to predict low-level control actions directly. For example, E2EParking [13] maps multi-view images to control outputs, but encodes goal semantics only implicitly within segmentation labels, which limits generalization. To enhance goal awareness, later methods [21, 26] explicitly represent parking slots—either by introducing slot-map queries fused with BEV features or by applying cross-attention between goal points and perception features. Despite these variations, previous end-to-end methods lack explicit attention mechanisms to ensure perception focuses on critical regions for planning and control. We address this with Control-Aided Attention (CAA), which is trained to highlight control-sensitive areas.
@@ II-B Deep Learning-based Motion Prediction
In our end-to-end parking setup, motion prediction is required to reliably track the ego vehicle’s state across multiple steps, ensuring that the planner operates on temporally consistent information. Prior end-to-end approaches such as E2EParking [13] employed segmentation-based tracking, but this design has been shown to fail in maintaining consistent predictions over multiple steps. Most traditional prediction methods are generally suitable for simple short-term tasks  [27, 28, 29], relying solely on the current state without modeling temporal dependencies. Recent temporal hybrid approaches combine multiple historical frames with physics-based constraints, employing architectures such as physics-informed RNN/LSTM [30, 31, 32], temporal CNN [33, 34], CNN–RNN hybrids [35, 36, 37], and attention mechanisms [38, 39, 40] to enhance long-horizon accuracy while maintaining physically feasible trajectories. Since our scenario is a static simulated environment, we adopt a lightweight CNN combined with a single-layer LSTM model. This design ensures model efficiency while leveraging temporal information to improve prediction generalization across different policies.
@@ II-C Trajectory as Planning Output
Trajectory-based methods are common in autonomous parking, as they explicitly generate a feasible path from the vehicle’s pose to a target slot. Recent works integrate learning modules to enhance robustness and comfort, with most models [41, 25, 21] directly predicting trajectories from observations for controller execution. Alternatively, TCP [7] uses predicted trajectories as intermediate guidance for control. Building on this idea, CAA-Policy employs trajectory prediction as an auxiliary supervisory signal, benefiting from its structural guidance without requiring online replanning.
@@ III Methodology
Figure 3: CAA-Policy consists of five main components: (1) Perception Backbone, (2) Feature Fusion, (3) Learnable motion prediction module, (4) Control-Aided Attention (CAA) module and (5) Prediction Heads. The prediction heads include control and waypoint prediction, while auxiliary heads for depth estimation and semantic segmentation are incorporated to improve feature learning. Notably, the Learnable motion prediction module is only used during inference, leveraging historical control and ego-motion information to reason about ego-vehicle dynamics and target position.
@@ III-A Problem Statement
  • Current vehicle state $s_{t}$, including scalar speed $v_{t}$, planar acceleration $\mathbf{a}_{t}=[a_{x},a_{y}]$, and heading (yaw) $\psi_{t}$.
  • Target parking goal $g_{t}=(x_{t}^{\text{target}},y_{t}^{\text{target}},\psi_{t}^{\text{target}})$, specified in the global coordinate system and transformed into the vehicle’s ego frame at the start of the episode.
  • Historical motion information, comprising four consecutive frames, along with control commands $c_{t-4\dots t-1}$, ego-motion, and vehicle states $s_{t-4\dots t-1}$, is used by the motion prediction module.
The system outputs the future positions and control commands:
  • The vehicle’s next-frame position $(x_{t+1},y_{t+1})$, which is also used to update the ego-frame target coordinates for subsequent steps.
  • Short-horizon control commands $c_{t+1...t+4}=(\text{steering}_{t+1...t+4},\text{throttle}_{t+1...t+4},\text{brake}_{t+1...t+4})$ for the next 4 time step.
  • Short-horizon 2D waypoints for the next 4 time steps, used to ensure smooth and consistent future trajectory: $\mathbf{w}_{t+1:t+4}=[(x_{t+1},y_{t+1}),\dots,(x_{t+4},y_{t+4})]$.
  • Auxiliary predictions for perception tasks, including depth and BEV semantics.
The objective is to train a supervised policy $\pi$ from expert trajectories $\mathcal{D}=\{\tau^{n}\}_{n=1}^{N}$, where N os the number of expert trajectories, and each $\tau^{n}=\{(X^{n}_{t},Y^{n}_{t})\}_{t=1}^{T_{n}}$. Each trajectory $\tau^{n}$ consists of $T_{n}$ time steps with inputs $X^{n}_{t}$ and corresponding expert outputs $Y^{n}_{t}$ at each step:
  • $X^{n}_{t}$: all inputs (multi-view images, vehicle state, target goal, historical motion).
  • $Y^{n}_{t}$: expert labels for all supervised tasks (control, next pose, waypoints, depth, BEV semantics).
  • $T_{n}$: length (number of time steps) of trajectory $\tau^{n}$.
@@
$\pi^{*}=\arg\min_{\pi}\frac{1}{N}\sum_{n=1}^{N}\frac{1}{T_{n}}\sum_{t=1}^{T_{n}}\mathcal{L}\left(Y_{t}^{n},\pi\left(X_{t}^{n}\right)\right),$
@@
Motion Prediction: Training vs. Inference. During training, the inputs do not require historical motion information, since ground-truth vehicle positions $\mathbf{w}_{t}$ and ego-target coordinates $\mathbf{g}_{0}$ are directly available and the motion prediction module is not used. In contrast, during inference, the module predicts these quantities from historical motion information to estimate the next-step vehicle state.
@@ III-B Network Architecture
CAA-Policy consists of five main stages: (1) Perception Backbone, (2) Feature Fusion, (3) Learnable motion prediction module, (4) Control-Aided Attention(CAA) module, and (5) Prediction Heads, as illustrated in Fig. 3. Auxiliary heads for depth and semantic segmentation are also included. Notions are explicitly shown in Tab. I
TABLE I: Summary of major notations for CAA-Policy, grouped by module. The Symbol column has a light gray background for easy distinction, while module headers have a very light blue background.
@@
$\mathbf{f}_{\text{BEV}}$ BEV feature from images
$\mathbf{f}_{\text{ego}}$ Encoded ego-motion
$\mathbf{f}_{\text{goal}}$ Encoded parking goal
$\mathbf{f}_{\text{fused}}$ Concatenated feature
@@
$c_{t}$ Control command (steering, throttle, brake)
$\mathbf{w}_{t+1:t+4}$ Short-horizon future 2D waypoints
$(x_{t+1},y_{t+1})$ Predicted next position
$h_{t}$ Dynamics LSTM hidden state
$\bm{\mu}_{t},\bm{\sigma}^{2}_{t}$ Predicted 2D displacement mean & variance
@@
AA Predicted BEV attention map (CAA)
$\mathbf{f}_{\text{CAA}}$ CAA-refined BEV feature
$\mathcal{L}_{\text{ctrl}},\mathcal{L}_{\text{wp}},\dots$ Task-specific losses
$\lambda_{\text{ctrl}},\lambda_{\text{wp}},\dots$ Loss weights

### v2

@@ Abstract
Precise parking requires an end-to-end system where perception adaptively provides policy-relevant details—especially in critical areas where fine control decisions are essential. End-to-end learning offers a unified framework by directly mapping sensor inputs to control actions, but existing approaches lack effective synergy between perception and control. Instead, we propose CAA-Policy, an end-to-end imitation learning system that allows control signal to guide the learning of visual attention via a novel Control-Aided Attention (CAA) mechanism. We train such an attention module in a self-supervised manner, using backpropagated gradients from the control outputs instead of from the training loss. This strategy encourages attention to focus on visual features that induce high variance in action outputs, rather than merely minimizing the training loss—a shift we demonstrate leads to a more robust and generalizable policy. To further strengthen the framework, CAA-Policy incorporates short-horizon waypoint prediction as an auxiliary task to improve temporal consistency of control outputs, a learnable motion prediction module to robustly track target slots over time, and a modified target tokenization scheme for more effective feature fusion. Extensive experiments in the CARLA simulator show that CAA-Policy consistently surpasses both the end-to-end learning baseline and the modular BEV segmentation + hybrid A* pipeline, achieving superior accuracy, robustness, and interpretability. Code and Collected Training datasets will be released. Code is released at https://github.com/ai4ce/CAAPolicy.
@@ I Introduction
Autonomous driving has traditionally been approached with modular pipelines that decompose the task into perception, prediction, and planning/control [grigorescu2020survey, chen2024end, chib2023recent]. While such designs provide interpretability and modularity [chen2015deepdriving, maddern20171, kong2015kinematic], they are prone to error accumulation across stages [wu2022trajectory], redundant computation [kong2015kinematic, zhao2012design], and reduced robustness under perturbations [wu2023adversarial],
Fig. 1LABEL:fig:teasing(a). End-to-end (E2E) learning offers a unified alternative by mapping sensory inputs directly to control outputs through a single neural network [hu2023planning, chitta2022transfuser, shao2023safety], Fig. 1LABEL:fig:teasing(b).
Existing end-to-end (E2E) autonomous parking methods [yang2024e2e] typically predict control commands directly from multi-view images. While these approaches unify perception and control, their synergy remains weak: perception modules are trained solely through gradients backpropagated from task losses, and thus often fail to focus on control-relevant cues such as parking slot boundaries, lane markings, and obstacle gaps. This limitation is particularly pronounced in attention mechanisms. As shown in Fig. 2(a–b), a baseline without control guidance scatters attention across irrelevant regions and exhibits temporal inconsistency. Moreover, even when incorporating an automatic learned attention module such as CBAM[woo2018cbam], the model still fails to consistently emphasize near-target areas. Without stronger coupling between perception and control, attention tends to drift toward spurious features, resulting in misaligned perception and unreliable policy outputs.
To address these limitations, we propose CAA-Policy, an end-to-end autonomous driving framework that tightly couples perception and control (Fig. 1LABEL:fig:teasing(c)). At its core, it introduces a self-supervised Control-Aided Attention (CAA) mechanism that learns attention from control-output gradients rather than the loss gradients, enabling the attention module to consistently focus on control-sensitive regions near the target slot (Fig. 2(c)). Driving goals are further encoded as structured, interpretable tokens for goal-aware control, and short-horizon waypoint prediction is added as an auxiliary task to improve temporal consistency. Together, these components yield more accurate and robust autonomous parking.
Figure 2: Decoder self-attention maps for target tracking at two consecutive frames ($t$ and t+1t{+}1). (a) Without CAA, self-attention is scattered over irrelevant areas; (b) With modified CBAM [woo2018cbam], an automatically learned attention module from task supervision (no auxiliary loss), attention is partially focused; (c) With CAA, self-attention is concentrated on the target and its vicinity; (d) Ground-truth segmentation (yellow: target, blue: obstacles) for reference.
@@
  • An end-to-end visual autonomous parking framework, CAA-Policy, that features a novel Control-Aided Attention (CAA) mechanism, which learns attention from control gradients on perception features, consistently focusing on control-sensitive areas.
  • We further improve performance by (1) introducing a learnable motion prediction to reliably track the target spot, (2) improving target tokenization to effectively encode the target input, and (3) adding a waypoint prediction auxiliary task to enhance planning stability.
  • Through extensive experiments in CARLA, we demonstrate that CAA-Policy  significantly outperforms both modular hybrid A* and previous end-to-end baselines in trajectory accuracy, success rate, and interpretability.
@@ II-A Autonomous parking system
Prior work on autonomous parking ranges from traditional modular pipelines to modern End-to-End (E2E) methods. Early systems [chirca2015autonomous, kim2025robust, zhang2025enhanced] typically combined occupancy mapping, hand-crafted planners (e.g., hybrid A* [dolgov2010path], RRT [lavalle1998rapidly]), and rule-based controllers, offering interpretability but limited adaptability. Recent advances largely adopt either modular learning approaches or E2E learning paradigms [al2025reinforcement, yang2024e2e, li2024parkinge2e].
Modular approaches follow the classic decomposition of perception, planning, and control, with each stage designed and tuned separately. Modular pipelines with hybrid A* planners [dolgov2008practical, kurzer2016path] face fundamental efficiency limitations that directly impact accuracy under practical trade-offs. Owing to the breadth-first-search nature of hybrid A*, planning becomes increasingly costly when the target is far, as the frontier space grows with each step. To mitigate the high computational burden, systems often reduce the frequency of re-planning with new perception inputs. However, this leads to blind execution based on outdated information, where occlusions near the target are frequently ignored, resulting in frequent failures. This shortcoming has motivated the subsequent preference for learning-based policies and end-to-end approaches.
End-to-end approaches provide a unified alternative to modular pipelines by directly mapping sensory inputs to control outputs, making them particularly suitable for structured tasks like autonomous parking. Within this paradigm, one line of work formulates parking as a trajectory prediction task, where CNN–LSTM or Transformer architectures are supervised with trajectory waypoints rather than direct control commands [shen2020parkpredict, shen2022parkpredict+, li2024parkinge2e]. Another line of work adopts imitation learning to predict low-level control actions directly. For example, E2EParking [yang2024e2e] maps multi-view images to control outputs, but encodes goal semantics only implicitly within segmentation labels, which limits generalization. To enhance goal awareness, later methods [li2024parkinge2e, fu2025parkformer] explicitly represent parking slots—either by introducing slot-map queries fused with BEV features or by applying cross-attention between goal points and perception features. Despite these variations, previous end-to-end methods lack explicit attention mechanisms to ensure perception focuses on critical regions for planning and control. We address this with Control-Aided Attention (CAA), which is trained to highlight control-sensitive areas.
@@ II-B Deep Learning-based Motion Prediction
In our end-to-end parking setup, motion prediction is required to reliably track the ego vehicle’s state across multiple steps, ensuring that the planner operates on temporally consistent information. Prior end-to-end approaches such as E2EParking [yang2024e2e] employed segmentation-based tracking, but this design has been shown to fail in maintaining consistent predictions over multiple steps. Most traditional prediction methods are generally suitable for simple short-term tasks  [prevost2007extended, yuan2021agentformer, huang2022survey], relying solely on the current state without modeling temporal dependencies. Recent temporal hybrid approaches combine multiple historical frames with physics-based constraints, employing architectures such as physics-informed RNN/LSTM [zyner2018recurrent, altche2017lstm, park2018sequence], temporal CNN [phan2020covernet, gilles2021home], CNN–RNN hybrids [deo2018convolutional, chandra2019traphic, xie2020motion], and attention mechanisms [kim2020multi, messaoud2020attention, giuliari2021transformer] to enhance long-horizon accuracy while maintaining physically feasible trajectories. Since our scenario is a static simulated environment, we adopt a lightweight CNN combined with a single-layer LSTM model. This design ensures model efficiency while leveraging temporal information to improve prediction generalization across different policies.
@@ II-C Trajectory as Planning Output
Trajectory-based methods are common in autonomous parking, as they explicitly generate a feasible path from the vehicle’s pose to a target slot. Recent works integrate learning modules to enhance robustness and comfort, with most models [du2025transparking, shen2022parkpredict+, li2024parkinge2e] directly predicting trajectories from observations for controller execution. Alternatively, TCP [wu2022trajectory] uses predicted trajectories as intermediate guidance for control. Building on this idea, CAA-Policy employs trajectory prediction as an auxiliary supervisory signal, benefiting from its structural guidance without requiring online replanning.
@@ III Methodology
Figure 3: CAA-Policy consists of five main components: (1) Perception Backbone, (2) Feature Fusion, (3) Learnable motion prediction module, (4) Control-Aided Attention (CAA) module and (5) Control and Waypoint Prediction. The network also incorporates auxiliary heads for depth estimation and semantic segmentation, following the design of E2EParking [yang2024e2e], to improve feature representation. Notably, the Learnable motion prediction module is only used during inference, leveraging historical control and vehicle state information to reason about ego-vehicle dynamics and target position.
@@ III-A Problem Statement
  • Current vehicle state $\mathbf{s}_{t}$, including scalar speed $v_{t}$, planar acceleration $\mathbf{a}_{t}$, and heading (yaw) $\psi_{t}$.
  • Target parking goal $\mathbf{g}_{t}=(x_{t}^{\text{target}},y_{t}^{\text{target}},\psi_{t}^{\text{target}})$, in the ego frame at time $t$.
  • Inputs to the motion prediction module: historical motion from four consecutive frames, control commands $\mathbf{c}_{t-4:t-1}$, and vehicle states $\mathbf{s}_{t-4:t-1}$.
The system outputs position, short-horizon control, and perception predictions:
  • Current-frame position $(x_{t},y_{t})$, predicted by the motion prediction module and used to update the ego-frame target coordinates for this step.
  • Short-horizon control commands for the next 4 steps, $\mathbf{c}_{t+1:t+4}=(\text{steering},\text{throttle},\text{brake})_{t+1:t+4}$.
  • Short-horizon 2D ego-centric waypoints for the next 4 time steps, used to ensure smooth and consistent future trajectory: $\mathbf{w}_{t+1:t+4}=[(x_{t+1},y_{t+1},\psi_{t+1}),\dots,(x_{t+4},y_{t+4},\psi_{t+4})]$.
  • Auxiliary outputs supporting perception tasks, including depth estimation and BEV semantics.
The objective is to train a supervised policy $\pi$ from expert trajectories $\mathcal{D}=\{\tau^{n}\}_{n=1}^{N}$, where $N$ is the number of expert trajectories. Each trajectory $\tau^{n}=\{(J^{n}_{t},K^{n}_{t})\}_{t=1}^{T_{n}}$ contains $T_{n}$ time steps:
  • $J^{n}_{t}$: multi-view images, vehicle state, target goal, and historical motion.
  • $K^{n}_{t}$: expert labels for all supervised tasks (control, next pose, waypoints, depth, BEV semantics).
@@
$\pi^{*}=\arg\min_{\pi}\frac{1}{N}\sum_{n=1}^{N}\frac{1}{T_{n}}\sum_{t=1}^{T_{n}}\mathcal{L}\left(K_{t}^{n},\pi\left(J_{t}^{n}\right)\right),$
@@ III-B Network Architecture
CAA-Policy consists of five main stages: (1) Perception Backbone, (2) Feature Fusion, (3) Learnable motion prediction module, (4) Control-Aided Attention(CAA) module, and (5) Control and Waypoint Prediction, as illustrated in Fig. 3. Auxiliary heads for depth and semantic segmentation are also included, following the design of E2EParking [yang2024e2e]. Notions are explicitly shown in Tab. I
TABLE I: Summary of major notations for CAA-Policy, grouped by module. The Symbol column has a light gray background for easy distinction, while module headers have a very light blue background. Notation conventions: continuous quantities are denoted by $\tilde{\cdot}$ (e.g., $\tilde{\mathbf{w}},\tilde{\mathbf{c}}$); ground-truth values are denoted by (e.g., $\mathbf{w}^{\ast},\mathbf{c}^{\ast}$); predicted distributions are denoted by $\hat{\cdot}$ (e.g., $\hat{\mathbf{w}},\hat{\mathbf{c}}$).
@@
$\mathbf{f}_{\text{BEV}}$ Tokenized BEV feature from images
$\mathbf{f}_{\text{ego}}$ Tokenized ego vehicle state vector
$\mathbf{f}_{\text{goal}}$ Tokenized parking goal vector
$\mathbf{f}_{\text{fused}}$ Concatenated Tokenized feature vector
@@
$\mathbf{c}_{t+1:t+4}$ Short-horizon control command vectors (steering, throttle, brake)
$\mathbf{w}_{t+1:t+4}$ Short-horizon future 2D waypoint vectors
$(x_{t},y_{t})$ Predicted current vehicle position (2D scalar coordinates)
$\mathbf{h}_{t}$ Dynamics LSTM hidden state vector
$\bm{\mu}_{t},\bm{\sigma}^{2}_{t}$ Predicted 2D displacement mean & variance vectors
@@
$\mathbf{A}$ Predicted BEV attention map (CAA)
$\nabla_{\mathbf{f}}^{\ast}$ Ground-truth control gradient w.r.t. fused feature
$\mathbf{f}_{\text{CAA}}$ CAA-refined BEV feature vector
$\mathcal{L}_{\text{c}},\mathcal{L}_{\text{w}},\dots$ Task-specific losses